S 1, S 2 ,S 3 …......... are the sum of infinite geometric series whose first terms are 1,2,3.......... and whose common ratios ½,1/3,1/4, respectivel,then S 1 2 +S 2 2 +...................S 10 2 is equal to – a> 485,b>495,c>500,d>505

S1,S2,S3 …......... are the sum of infinite geometric series whose first terms are 1,2,3.......... and whose common ratios ½,1/3,1/4, respectivel,then S1+S22+...................S10 is  equal to – 
a> 485,b>495,c>500,d>505

Grade:

3 Answers

manasa gorthy
9 years ago
S1= 1/(1-1/2)=2                                                                                                                             S2= 2/(1-1/3)=3   similarly S10=11                                                                                                   S12+S22+...................S102=22+32+..............112=505              
Vamsi Krishna Velpuri
9 years ago
S1= 1/(1-1/2)=2                                                                                                                             S2= 2/(1-1/3)=3   similarly S10=11                                                                                                   S12+S22+...................S102=22+32+..............112
=(∑ n2) -1 where n=1 to 11
=>answer is 505-1=504
Vamsi Krishna Velpuri
9 years ago
sorry it is 506-1=505